A circular loop of wire of radius 0.25 cm is expected to an external magnetic field that is perpendicular to the plane of the loop. For a time, the magnetic field decreases at a rate of 0.0035 T/s. What is the magnitude of the induced electromagnetic force in the loop?



Show step by step solution:



Given:



Solution:

1
Expert's answer
2022-05-22T14:39:27-0400

"F=SB\cos\alpha=\pi r^2 B=6.9\cdot 10^{-8}~N."